OpenStudy (anonymous):

7. Short Answer. The length of the side of a rectangular desk is given by 3x - 2y cm. The length of the front is given by 3x + 2y cm. What is an algebraic expression for the area of the top of the rectangular desk? Include correct units as part of your answer. (Points : 3)

OpenStudy (anonymous):

Length of side= width = 3x -2y cm Length of front = length = 3x +2y cm Area = length * width = (3x -2y)*(3x +2y) sq. cm Area = (3x)^2 -(2y)^2 sq. cm Area = 9x^2 - 4y^2 sq. cm